# Gas Transfer Between the Inner 3-kpc Disk and the Galactic Central Molecular Zone

### v1

@@ 4 Large-scale Tilted Gas Flows and The Twisted Structure in the CMZ
The dynamical processes in the inner Galaxy cause gas flows to exchange mass and angular momentum with their surroundings, enabling the CMZ to gain fresh gas from the Galactic disk and the recycling gas from the overshoot effect. Given the long-term accretion process and the remarkable transient accretion effects caused by overshooting, the non-uniform accumulation of gas (e.g., the instantaneous accretion rate of about 5–10 ) can significantly feed and regulate star formation in the CMZ. Due to the highly distorted and inhomogeneous nature of gas inflows, the accretion process onto the CMZ, and even onto the SMBH, are likely highly non-uniform. This could result in the ring of the CMZ and jets of the SMBH exhibiting strong precession and intermittent variability. Furthermore, feedbacks from star-formation in the CMZ and the SMBH (e.g., Emsellem et al., 2015; Kruijssen et al., 2015; Krumholz et al., 2017, 2018; Armillotta et al., 2019; Anderson et al., 2020; Sormani et al., 2020; Tress et al., 2020; Martínez-Arranz et al., 2024; Takekawa et al., 2024) contribute to reducing the angular momentum of gas in the CMZ, enabling a fraction of lower-angular-momentum gas to migrate further towards the region closer to the GC and thereby increasing the potential activities of the SMBH.
@@
In simple terms, the gas and star-forming environment in the CMZ can be regulated by both the external gas input through the bar channel (this work) and the various feedback therein (e.g., Heywood et al., 2019; Ponti et al., 2021; Mackey et al., 2024; Nogueras-Lara, 2024; Nonhebel et al., 2024; Sarkar, 2024), thus achieving a delicate balance between gas fuel supply and star formation in the extreme environments. Under the control of non-axisymmetric gravitational potential of the Galactic bar, the continuous accumulations of gas with VARYING inflow rates, along with frequent interactions between gas flows and star formation, can exist simultaneously within a few hundreds parsecs of the GC over a long timescale. This is what distinguishes the CMZ from other star-forming regions on the Galactic disk.
@@ 5 Summary
we investigate the nature of the approaching and receding gas flows towards
@@ 1. The MWISP CO data show that a large-scale gas structure, characterized by
towards the observer (Figure 1). Based on the observations,
@@
2. The overshoot clouds are decelerating (G7), re-accelerating (G5), and
@@
is constantly changing its state of motion and spiraling towards the CMZ
@@
the overshoot effect, is about two times of the fresh gas accreted directly
@@
3. The cloud-cloud collision features between the overshoot gas from the
far dust lane and the gas inflows towards the CMZ are clearly evident
in G3 and G5 regions (Figure 2; see details in Section 2).
@@ 3. The cloud-cloud collision features between the overshoot gas from the
shearing/non-relaxed gas features relative to the bar, frequent collisions
@@ 4. As one of the most important gas components, molecular gas undergoes
non-negligible DISSIPATION processes due to strong turbulence. During
@@ 5. The overshoot flows exhibit a more pronounced tilted feature compared
the tilted gas flows on kpc scales and the local warped gas structure
dominated by the strong gravitational potential of the nuclear stellar disc.
The non-axisymmetric gravitational potential of the Galactic bar
@@
6. The gas inflows rate towards the CMZ is 1.1 yr-1
@@
and the overshoot gas from the far dust lane. The average gas inflow rate towards
@@ 7. The MWISP survey with the high-sensitivity and high-velocity
feedbacks within a few hundred parsecs of the GC, jointly shape the gas
@@
The holistic structure and large-sample statistical properties revealed by the observations are important for our understanding of the gas dynamics towards the center of galaxies. Moreover, many simulation studies have been done to explain how gas structure form within a bar and how they depend on various bar parameters (e.g., the bar strength, the effective sound speed, the pattern speed of the bar, the inclination angle of the bar, the magnetic field, and the various gravitational potential, etc.; see Lindblad et al., 1996; Kim et al., 2012a, b; Kim & Stone, 2012; Sormani et al., 2024; Feng et al., 2024; Tress et al., 2024). The comprehensive analysis of large-scale survey data, combined with the comparison of observational features and model simulations, is essential for accurately understanding the astrophysical phenomena and fundamental physical processes that dominate galaxy evolution.

### v2

@@ 4 Large-scale Tilted Gas Flows and the Twisted Structure in the CMZ
The dynamical processes in the inner Galaxy cause gas flows to exchange mass and angular momentum with their surroundings, enabling the CMZ to gain fresh gas from the Galactic disk and the recycled gas from the overshoot effect. Given the long-term accretion process and the remarkable transient accretion effects caused by overshooting, the nonuniform accumulation of gas (e.g., the instantaneous accretion rate of about 5–10 ) can significantly feed and regulate star formation in the CMZ. Due to the highly distorted and inhomogeneous nature of gas inflows, the accretion process onto the CMZ, and even onto the SMBH, is likely highly nonuniform. This could result in the ring of the CMZ and jets of the SMBH exhibiting strong precession and intermittent variability. Furthermore, feedback from star-formation in the CMZ and the SMBH (e.g., Emsellem et al., 2015; Kruijssen et al., 2015; Krumholz et al., 2017, 2018; Armillotta et al., 2019; Anderson et al., 2020; Sormani et al., 2020; Tress et al., 2020; Martínez-Arranz et al., 2024; Takekawa et al., 2024) contributes to reducing the angular momentum of gas in the CMZ, enabling a fraction of lower angular momentum gas to migrate further toward the region closer to the GC and thereby increasing the potential activities of the SMBH.
@@
In simple terms, the gas and star-forming environment in the CMZ can be regulated by both the external gas input through the bar channel (this work) and the various feedback therein (e.g., Heywood et al., 2019; Ponti et al., 2021; Mackey et al., 2024; Nogueras-Lara, 2024; Nonhebel et al., 2024; Sarkar, 2024), thus achieving a delicate balance between gas fuel supply and star formation in the extreme environments. Under the control of nonaxisymmetric gravitational potential of the Galactic bar, the continuous accumulations of gas with VARYING inflow rates, along with frequent interactions between gas flows and star formation, can exist simultaneously within a few hundred parsecs of the GC over a long timescale. This is what distinguishes the CMZ from other star-forming regions on the Galactic disk.
@@ 5 Summary
we investigate the nature of the approaching and receding gas flows toward
@@ 1. The MWISP CO data show that a large-scale gas structure, characterized by
toward the observer (Figure 1). Based on the observations,
@@
2. The overshoot clouds are decelerating (G7), reaccelerating (G5), and
@@
is constantly changing its state of motion and spiraling toward the CMZ
@@
the overshoot effect, is about 2 times that of the fresh gas accreted directly
@@
3. The cloud–cloud collision features between the overshoot gas from the
far dust lane and the gas inflows toward the CMZ are clearly evident
in the G3 and G5 regions (Figure 2; see details in Section 2).
@@ 3. The cloud–cloud collision features between the overshoot gas from the
shearing/nonrelaxed gas features relative to the bar, frequent collisions
@@ 4. As one of the most important gas components, molecular gas undergoes
nonnegligible DISSIPATION processes due to strong turbulence. During
@@ 5. The overshoot flows exhibit a more pronounced tilted feature compared
the tilted gas flows on kiloparsec scales and the local warped gas structure
dominated by the strong gravitational potential of the nuclear stellar disk.
The nonaxisymmetric gravitational potential of the Galactic bar
@@
6. The gas inflow rate toward the CMZ is 1.1 yr-1
@@
and the overshoot gas from the far dust lane. The average gas inflow rate toward
@@ 7. The MWISP survey with the high-sensitivity and high-velocity
feedback within a few hundred parsecs of the GC, jointly shape the gas
@@
The holistic structure and large-sample statistical properties revealed by the observations are important for our understanding of the gas dynamics toward the center of galaxies. Moreover, many simulation studies have been done to explain how gas structures form within a bar and how they depend on various bar parameters (e.g., the bar strength, the effective sound speed, the pattern speed of the bar, the inclination angle of the bar, the magnetic field, the various gravitational potentials, etc.; see Lindblad et al., 1996; Kim et al., 2012a, b; Kim & Stone, 2012; Feng et al., 2024; Sormani et al., 2024; Tress et al., 2024). The comprehensive analysis of large-scale survey data, combined with the comparison of observational features and model simulations, is essential for accurately understanding the astrophysical phenomena and fundamental physical processes that dominate galaxy evolution.